OpenStudy (anonymous):

A garden hose is used to fill a large metal container which can hold 21.0 L of liquid. If the radius of the garden hose's nozzle is 1.1 cm, and the speed of the water at this point is 273.0 cm/s, how long (in seconds) would it take to fill the container with water, assuming that none of it splashes out?

OpenStudy (anonymous):

We can model the rate of volume flow as\[\huge Q= \int\limits\int\limits_A v \cdot dA \implies Q=v_\text{avg}A\]Where: Q = rate of flow v = viscous flow (average velocity) A = cross sectional area The time it takes to fill a certain volume while observing the rate of volume flow is\[\huge t=\frac{V}{Q}\]Where: t = time to fill the volume V = volume Q = rate of flow

OpenStudy (michele_laino):

correct!! @CShrix

OpenStudy (anonymous):

Yay :D